# Multi-LF: A Unified Continuous Learning Framework for Real-Time DDoS Detection in Multi-Environment Networks

### v1

@@ IV-C Results in Real-Time Environment
Table IX provides an evaluation of Scenario 1, where the M1 model is used without continuous learning for attack detection. KNN demonstrates the highest accuracy (0.688) but at the cost of substantial resource usage, including a large MS (90,236 KB), high memory consumption (187.741 MB), and significant CPU utilization (58.34%). In contrast, models like LR show a much smaller size (2 KB), lower MU (122.3 MB), and minimal CPU impact (45.61%) with a faster prediction time (0.0007 seconds), though with reduced accuracy (0.536). This highlights the trade-off between model complexity and computational efficiency, making KNN suitable for accuracy-critical scenarios, while lighter models like LR are better for resource-constrained environments.
TABLE IX: Scenario 1: M1 Without Continuous Learning
@@
MU (MB) 122.467 121.571 187.741 122.3 112.81
Scenario 2 evaluates the M1 model with continuous learning capabilities, showcasing improvements in both efficiency and performance as shown in Table X. Among the models tested, Perceptron achieves the highest accuracy (0.704) while maintaining a minimal MS (3 KB) and low prediction time (0.0007 seconds). This demonstrates the model’s ability to quickly adapt to new data without a significant increase in computational costs. Other models like MNB and SGDC also display competitive accuracies (0.594 and 0.581, respectively) and maintain a low memory footprint (around 117-120 MB). These results highlight the effectiveness of continuous learning in maintaining high accuracy while optimizing resource usage for real-time scenarios as it achieved 0.704 high accuracy as compared to Scenario 1 which has 0.688 accuracy.
TABLE X: Scenario 2: M1 With Continuous Learning
Model BNB MNB Perceptron SGDC
MS (KB) 4 4 3 3
@@
MU (MB) 110.248 117.812 124.405 120.932
In Scenario 3, the performance of M2 is evaluated without continuous learning across multiple ML models, as shown in Table XI. The results indicate a significant variation in accuracy, prediction time, CPU utilization, and MU among the models. MLP achieves the highest accuracy (0.885), demonstrating strong prediction capabilities, but at the cost of increased CPU utilization (57.14%) and MU (140.119 MB). On the other hand, LR provides a balanced performance with good accuracy (0.818) and low computational costs, making it also a suitable choice for applications with limited resources. RF model, while offering moderate accuracy (0.738), has the largest MS (33,009 KB) and the highest memory consumption (267.679 MB), indicating a trade-off between performance and resource efficiency. Models like BNB and GNB, with accuracies of 0.725 and 0.721, respectively, exhibit low prediction time and MU, but their high CPU utilization makes them less practical for real-time applications. Compared to Scenario 2, where continuous learning was employed, Scenario 3 shows that M2 without continuous learning achieves higher accuracy in some models, such as MLP, but with increased computational overhead, indicating the benefits of continuous learning for resource optimization and real-time applicability. Considering these factors, we utilize M1 with continuous learning at the first level and then deploy the best-performing model from M2 at the second level in Multi-LF.
TABLE XI: Scenario 3: M2 Without Continuous Learning
Model MS (KB) Accuracy PT (S) CPU (%) MU (MB)
ADA 57 0.723 0.085 39.91 96.179
@@
SGDC 3 0.830 0.002 46.00 137.654
Table XII shows the performance comparison of four models, BNB, MNB, Perceptron, and SGDC, under Scenario 4, where both M1 and M2 models are deployed in Multi-LF without human involvement. The first set of results evaluates the models with MLP as the secondary model (M2). The Perceptron model achieves the highest accuracy at 0.999, but at the cost of significantly higher CPU utilization (23.22%) and MU (2.623 MB), indicating its computational intensity. In contrast, BNB maintains high accuracy at 0.975 with a much lower CPU usage (2.582%) and MU (1.513 MB), suggesting a good trade-off between performance and resource efficiency.
The second set of results involves the same models but with RF as the M2. The accuracy of all models drops compared to the MLP scenario, with the highest accuracy achieved by the Perceptron model (0.983). However, the computational overhead is significantly reduced, as evidenced by lower CPU utilization (3.221%) and MU (0.956 MB). This suggests that using RF as the secondary model is less resource-intensive, but it comes at the cost of reduced prediction performance. The results highlight the trade-offs between model performance and resource efficiency when using different secondary models (MLP and RF) within Multi-LF. The Perceptron model consistently achieves high accuracy but at a higher computational cost, while BNB and other models offer a balanced approach between accuracy and resource usage.
TABLE XII: Scenario 4: M1 and M2 Under Multi-LF Without Human Involvement
MLP as M2 & Without Human Check
Matrix BNB MNB Preceptron SGDC
Accuracy 0.975 0.331 0.999 0.371
@@
RF as M2 & Without Human Check
Accuracy 0.941 0.330 0.983 0.186
@@
In Scenario 5, M1 and M2 models were evaluated under the Multi-LF with human involvement. The models used were BNB, MNB, Perceptron, and SGDC, with MLP and RF as the secondary models (M2), same as Scenario 4. This scenario demonstrated the impact of incorporating human oversight on model performance. When MLP was used as the secondary model, the Perceptron achieved the highest accuracy of 0.999, but it required significant CPU utilization (17.84%) and MU (35.48 MB). The prediction time of 0.866 seconds indicates that while human involvement boosts the model’s accuracy, it also increases computational costs. In contrast, the BNB model showed good performance with an accuracy of 0.979, but its MU (56.53 MB) remained notably high compared to other models. When RF was used as the secondary model, Perceptron again demonstrated high accuracy (0.983) but with a reduced CPU utilization of 10.05% and lower MU (3.632 MB). Prediction time was also reduced to 0.434 seconds, indicating that RF as the secondary model results in a more resource-efficient configuration while maintaining high accuracy. BNB’s performance with RF was similar to its performance with MLP in terms of accuracy, but it achieved lower CPU utilization (2.564%) and higher MU (72.69 MB).
Comparing Scenario 5 to Scenario 4 Multi-LF, we observe that human intervention slightly increases computational resource usage but improves model accuracy, demonstrating the value of human oversight. Furthermore, Scenario 5 outperformed Scenario 1 and Scenario 2 (M1 configurations) in accuracy across most models, validating the robustness of the proposed human-involved framework. Finally, compared to Scenario 3 (M2 without continuous learning), Scenario 5 exhibited a more balanced trade-off between accuracy and computational efficiency, especially when using RF as the secondary model, highlighting the advantage of human intervention in refining the model’s performance.
TABLE XIII: Scenario 5: M1 and M2 Under Multi-LF With Human Involvement
MLP as M2 & Human Check
Matrix BNB MNB Preceptron SGDC
Accuracy 0.979 0.331 0.999 0.371
@@
(a) Performance of M1 (BNB) and M2 (RF) Models without Human Intervention
(b) Performance of M1 (BNB) and M2 (MLP) Models without Human Intervention
(c) Performance of M1 (BNB) and M2 (RF) Models with Human Intervention
(d) Performance of M1 (BNB) and M2 (MLP) Models with Human Intervention
(e) Performance of M1 (Perceptron) and M2 (RF) Models without Human Intervention
(f) Performance of M1 (Perceptron) and M2 (MLP) Models without Human Intervention
(g) Performance of M1 (Perceptron) and M2 (RF) Models with Human Intervention
(h) Performance of M1 (Perceptron) and M2 (MLP) Models with Human Intervention
Figure 7: Scenario 4 and Scenario 5: Best performing models’ results per batch
@@
IV-D Comparison With Benchmark Studies
Many studies validate their MTDS approaches in offline settings, often achieving significant results. However, these systems frequently experience performance degradation when deployed in real-time environments due to the dynamic nature of network traffic and the emergence of new attack patterns. The datasets used to train these models are typically static, which limits their effectiveness in real-world applications.
@@ IV-D Comparison With Benchmark Studies
There are no existing studies that evaluate their work on real-time online M-En traffic. Therefore, for a fair comparison, we implemented existing M-En studies on our newly collected real-time M-En dataset, deploying them according to the methodologies outlined in their respective published work. Since most existing studies conducted only offline testing, we replicated their models in the same offline setting to maintain consistency in evaluation. For instance, Rustam et al. [47] deployed their approach on an M-En dataset using the synthetic data augmentation technique (S-DATE). While the performance appeared strong after applying the S-DATE data balancing method, the results indicated potential data leakage issues, which could explain the observed high accuracy. Similarly, other studies [14, 10] proposed optimization-based approaches to handle diverse traffic patterns in M-En networks, employing Particle Swarm Optimization (PSO) and Moth Flame Optimization (MFO), respectively. These studies introduced PSO-diverse self-ensemble model (PSO-D-SEM) and a Fully Automated Malicious Traffic Detection System (FAMTDS). Another study [70] utilized a dual-data trained LightGBM (DDT-LightGBM) model, achieving significant results in the M-En network within an offline testing framework. Zukaib et al. [49] integrated Federated Learning and Meta-learning to propose the Meta-Fed IDS, which was tested on the M-En dataset. However, their evaluation was limited to offline settings and did not incorporate real-time data collection or adaptation. In contrast, our study conducted both online and offline testing and was built on a real-time M-En dataset.
In Table XIV, we compare our approach, Multi-LF, to relevant M-En benchmark studies. Most existing approaches [47, 14, 10, 70, 49] use static datasets and evaluate performance only in offline scenarios, reporting accuracy values between 0.94 and 0.991. While these systems can be effective in controlled settings, they often lack the mechanisms to continuously learn from new traffic patterns or adapt to real-time fluctuations in M-En networks.
@@
In contrast, Multi-LF is the only method evaluated in both offline and online modes, achieving near-optimal accuracy scores of 1.00 (offline) and 0.999 (online). These results highlight Multi-LF’s resilience to dynamic traffic behavior. By continuously retraining on fresh data and utilizing multi-level validation checks, Multi-LF maintains high accuracy even under unpredictable network conditions, offering a robust solution for real-time MTDS in M-En environments.
TABLE XIV: Comparison With Existing Studies
@@
[47] 2023 ETC, S-DATE $\checkmark$ $\times$ 0.986
[14] 2024 PSO-D-SEM $\checkmark$ $\times$ 0.978
[10] 2024 FAMTDS $\checkmark$ $\times$ 0.991
[70] 2024 DDT-LightGBM $\checkmark$ $\times$ 0.98
[49] 2024 Meta-Fed IDS $\checkmark$ $\times$ 0.94
- 2025 Multi-LF $\checkmark$ $\checkmark$ 1.00, 0.999
V Discussion
This study collected a dataset in the M-En environment and proposed the Multi-LF framework for attack detection, testing it in both offline and real-time scenarios. We deployed Multi-LF to reduce computational costs and enhance accuracy over time. The computational cost is reduced because the initial traffic is evaluated by the lightweight M1 model, which has a faster prediction time, as shown in Table VIII. Most of the traffic is filtered by M1, and only a small portion is forwarded to M2 and, if necessary, to a human expert. Additionally, the approach’s performance improved by incorporating an extra layer of security using M2 and human involvement. This is reflected in the improvement in accuracy from 0.70 to 0.99, as shown in Figure 8.
Figure 8: Accuracy comparison with difference scenarios
The improvement in model performance was achieved through continuous training with new data, allowing the model to efficiently adapt to new traffic patterns as well as through our feature engineering approach. The use of statistical features in our methodology significantly contributed to the model’s success, as shown in Figure 9. Statistical features such as ConnectionErrors, DstPortEntropy, MostFreqPayloadSize, SourceEntropy, MostFreqPacketSizeFreq, FlowRate, PacketCount, PacketSizeVar, and AvgPayloadSize played a major role in detecting DDoS attacks. For example, a sudden increase in packet count from a specific IP within a time window can indicate an attack. These statistical features, combined with general features, generated highly correlated inputs, as shown in Figure 3, thus improving the performance of the framework.

### v2

@@ 4.3 Results in Real-Time Environment
Table 9 provides an evaluation of Scenario 1, where the M1 model is used without continuous learning for attack detection. KNN demonstrates the highest accuracy of 0.688 but at the cost of substantial resource usage, including a large MS 90,236 KB, high memory consumption 187.741 MB, and significant CPU utilization 58.34%. In contrast, models like LR exhibit a much smaller size of 2 KB, a lower MU of 122.3 MB, and a minimal CPU impact of 45.61% with a faster prediction time of 0.0007 seconds, with reduced accuracy of 0.536. These results highlight the trade-off between model complexity and computational efficiency, making KNN suitable for accuracy-critical scenarios, while lighter models, such as LR, are better suited for resource-constrained environments.
Table 9: Scenario 1: Baseline M1 Without Continuous Learning
@@
[…]
Scenario 2 evaluates the M1 model with continuous learning capabilities, showcasing improvements in both efficiency and performance as shown in Table 10. Among the models tested, PER achieves the highest accuracy of 0.704 while maintaining a minimal MS of 3 KB and a low prediction time of 0.0007 seconds. These results demonstrate the model’s ability to quickly adapt to new data without a significant increase in computational costs. Other models, such as MNB and SGDC, also display competitive accuracies of 0.594 and 0.581, respectively, while maintaining a low memory footprint of around 117-120 MB. These results highlight the effectiveness of continuous learning in maintaining high accuracy while optimizing resource usage for real-time scenarios, as it achieves an accuracy of 0.704 compared to Scenario 1, which has an accuracy of 0.688.
Table 10: Scenario 2: Baseline M1 With Continuous Learning
Model BNB MNB PER SGDC
[…]
@@
[…]
In Scenario 3, the performance of M2 is evaluated without continuous learning across multiple machine learning models, as shown in Table 11. The results show notable variation in accuracy, prediction time, CPU utilization, and memory usage. MLP achieves the highest accuracy of 0.885 but requires high CPU utilization of 57.14% and memory usage of 140.12 MB. LR offers balanced performance with an accuracy of 0.818 and low computational cost, making it suitable for resource-constrained environments. RF attains moderate accuracy of 0.738 but consumes the most resources, with an MS of 33,009 KB and MU of 267.68 MB, highlighting a trade-off between accuracy and efficiency. BNB and GNB achieve accuracies of 0.725 and 0.721 respectively, maintaining low prediction times and memory use, though their high CPU utilization limits real-time applicability. Compared to Scenario 2, where continuous learning was employed, Scenario 3 shows that M2 can achieve higher accuracy with some models, such as MLP, but at the cost of increased computational overhead. This emphasizes the importance of continuous learning for optimizing resource usage and ensuring real-time performance. Consequently, M1 with continuous learning is applied at the first level, and the best M2 model is deployed at the second level in the Multi-LF.
Table 11: Scenario 3: Baseline M2 Performance Comparison
Model MS (KB) Acc PT (S) CPU (%) MU (MB)
[…]
@@
[…]
Table 12 shows the performance comparison of four models, BNB, MNB, PER, and SGDC, under Scenario 4, where both M1 and M2 models are deployed in Multi-LF without human involvement. The first set of results evaluates the models with MLP as the M2 because of its significant performance. The PER model achieves the highest accuracy at 0.999, but at the cost of significantly higher CPU utilization, 23.22%, and MU 2.623 MB, indicating its computational intensity. In contrast, BNB maintains high accuracy at 0.975 with a much lower CPU usage of 2.582% and MU 1.513 MB, suggesting a good trade-off between performance and resource efficiency.
The second set of results involves the same models but with RF as the M2. The accuracy of all models drops compared to the MLP scenario, with the highest accuracy of 0.983, achieved by the PER model. However, the computational overhead is significantly reduced, as evidenced by lower CPU utilization, 3.221%, and MU 0.956 MB. These results suggest that using RF as the secondary model is less resource-intensive, but it comes at the cost of reduced prediction performance. The results highlight the performance, efficiency trade-offs between using MLP and RF as M2 models in Multi-LF. The PER model consistently achieves high accuracy but at a higher computational cost, while BNB and other models offer a balanced approach between accuracy and resource usage.
Table 12: Scenario 4: Baseline M1 and M2 Under Multi-LF Without Human-in-the-Loop
MLP → M2 & Without Human-in-the-Loop
Matrix BNB MNB PER SGDC
[…]
@@
RF → M2 & Without Human-in-the-Loop
Matrix BNB MNB PER SGDC
[…]
@@
In Scenario 5, M1 and M2 models were evaluated under the Multi-LF with human involvement. We used BNB, MNB, PER, and SGDC as M1 with MLP and RF as the M2, same as Scenario 4. This scenario demonstrated the impact of incorporating human oversight on model performance. When MLP was used as the M2, the PER achieved the highest accuracy of 0.999, but it required significant CPU utilization of 17.84% and MU of 35.48 MB. The prediction time of 0.866 seconds indicates that while human involvement boosts the model’s accuracy, it also increases computational costs. In contrast, the BNB model showed good performance with an accuracy of 0.979, but its MU of 56.53 MB remained notably high compared to other models. When RF was used as the M2, PER again demonstrated high accuracy of 0.983, but with a reduced CPU utilization of 10.05% and a lower MU of 3.632 MB. Prediction time was also reduced to 0.434 seconds, indicating that RF as the M2 results in a more resource-efficient configuration while maintaining high accuracy. BNB’s performance with RF was similar to its performance with MLP in terms of accuracy, but it achieved lower CPU utilization, 2.564%, and higher MU of 72.69 MB. Comparing Scenario 5 to Scenario 4 under Multi-LF, we observe that human intervention slightly increases computational resource usage but improves model accuracy, demonstrating the value of human oversight. Furthermore, Scenario 5 outperformed Scenario 1 and Scenario 2 (M1 configurations) in accuracy across most models, validating the robustness of the proposed human-involved framework. Finally, compared to Scenario 3 (M2 without continuous learning), Scenario 5 exhibited a more balanced trade-off between accuracy and computational efficiency, especially when using RF as the M2, highlighting the advantage of human intervention in refining the model’s performance.
Table 13: Scenario 5: Baseline M1 and M2 Under Multi-LF With Human-in-the-Loop
MLP → M2 & Human-in-the-Loop
Matrix BNB MNB PER SGDC
[…]
@@
Human Intervention Feasibility: In our experimental setup, approximately 111,644 samples are processed across 15 batches. Human intervention is required for only 109 samples, resulting in a human effort percentage of just 0.0976%. This behavior shows that the approach is highly feasible for real-time deployment, as the burden on human input remains minimal, as shown in Table 14.
Table 14: Human Workload Under Multi-LF  Using Baseline M1 & M2
Metric M1 → M2 → Human
[…]
M1 CPP 111,153
[…]
M2 CPP 382
[…]
HE 0.0976%
  • TRP– Total Received Packets; CPP– Correctly Predicted Packets; HIR– Human Intervention Required; HE– Human Efforts
@@
(a) M1 (BNB) & M2 (RF) without Human-in-the-Loop
(b) M1 (BNB) & M2 (MLP) without Human-in-the-Loop
(c) M1 (BNB) & M2 (RF) with Human-in-the-Loop
(d) M1 (BNB) & M2 (MLP) with Human-in-the-Loop
(e) M1 (PER) & M2 (RF) without Human-in-the-Loop
(f) M1 (PER) & M2 (MLP) without Human-in-the-Loop
(g) M1 (PER) & M2 (RF) with Human-in-the-Loop
(h) M1 (PER) & M2 (MLP) with Human-in-the-Loop
Figure 8: Per-Batch Results of Best Models For Scenarios 4 and 5
4.4 Proposed DL Models Results Under Multi-LF Framework
In addition to traditional machine learning models, we implemented several deep learning architectures in our proposed Multi-LF  to enhance malicious traffic detection performance. These include T1, T2, L1, and L2, where T1 and L1 represent lightweight models (used as M1) and T2 and L2 represent more complex variants (used as M2). Deep learning models showed significant performance improvements compared to traditional baselines, largely due to their ability to capture long-term dependencies and complex patterns in the data. The lightweight variants were specifically designed to maximize throughput while maintaining high accuracy. Table 15 shows DL models’ performance achieved nearly identical score values, exceeding 0.999, indicating robust and reliable performance. The lightweight models (T1 and L1) show higher throughput, especially T1, which achieved a throughput of 66,130 packets/sec, significantly higher than L1 (6,157 packets/sec) and the complex models T2 and L2. Notably, T1 had a slightly higher error rate of 0.0651, compared to L1, which is 0.0612, but the difference is minimal and offset by its efficiency advantage.
Table 15: Results Using Proposed DL-based Models
Model Acc Pre Rec F1 Error Rate Throughput
[…]
T2 0.9994 0.9994 0.9994 0.9994 0.0612 65595
[…]
L2 0.9994 0.9995 0.9994 0.9994 0.0611 4263
Table 16 further summarizes the performance of our Multi-LF  under human-in-the-loop settings. Both the T1 $\rightarrow$ T2 and L1 $\rightarrow$ L2 configurations correctly handled nearly all incoming samples at the M1 stage (115,789 and 116,669 correctly predicted packets, respectively). Only a few packets required M2-level processing (6 for T1, 4 for L1), and even fewer required human intervention (3 for T1, none for L1). The human effort (HE) remained minimal in both cases, just 0.0026% for the Transformer-based path and 0.0000% for the LSTM-based path. Due to the higher throughput and lower error rate, we selected the Transformer-based architecture for deployment over the LSTM-based one. The minimal involvement required from human experts proves that our approach is both feasible and effective for real-time scenarios.
Table 16: Human Workload Under Multi-LF  Using DL-based M1 & M2
Metric T1 → T2 → Human L1 → L2 → Human
[…]
M1 CPP 115,789 116,669
[…]
M2 CPP 3 4
[…]
HE 0.0026 % 0.0000 %
  • TRP– Total Received Packets; CPP– Correctly Predicted Packets; HIR— Human Intervention Required; HE— Human Efforts
Further, to demonstrate the significance and sustainability of our proposed approach under the Multi-LF framework, we measured the environmental impact and computational cost using the CodeCarbon [^4] https://codecarbon.io/ library. Table 17 presents the carbon emissions and energy consumption metrics for each deep learning model variant. The results show that the lightweight models (T1 and L1) are notably more energy-efficient and environmentally friendly than their complex counterparts (T2 and L2). For instance, the T1 model consumed only 0.0242 kWh of energy and emitted just 0.0070 kg of CO2, while the L2 model, being the most complex, used 0.1074 kWh and emitted 0.0312 kg of CO2.
Table 17: Environmental Impact Statistics of DL-based Models
Model CO2 CPU_E RAM_E Total_E CPU_P
[…]
T2 0.0134 0.0314 0.0148 0.0462 42.5
[…]
L2 0.0312 0.0730 0.0343 0.1074 42.5
  • CO2– CO2 Emission (KG); CPU_E– CPU Energy (kWh); RAM_E– RAM Energy (kWh); Total_E–Total Energy (kWh); CPU_P– CPU Power (W)
4.5 Results using Other Benchmark Datasets
This section presents the performance of the proposed Multi-LF framework using two benchmark datasets, M-En Dataset-1 [rustam2024malicious] and M-En Dataset-2 [rustam2024famtds], which are synthetically created by combining well-known intrusion detection datasets. Dataset-1 is a combination of IoTID-20, UNSW-NB15, and InSDN, while Dataset-2 consists of IoTID-20 and UNSW-NB15. The results show that the M1 correctly classifies the majority of packets, while the M2 effectively handles low-confidence cases passed from M1, as shown in Table 18. Only a small fraction of packets require human intervention, with human effort reduced to as low as 0.29% in Dataset-2. These significant results show the effectiveness of the proposed Multi-LF in maintaining high accuracy while substantially reducing human workload across different benchmark datasets.
Table 18: Proposed Multi-LF  Using Benchmark Datasets
Metric Dataset-1 [rustam2024malicious] Dataset-2 [rustam2024famtds]
[…]
M1 CPP 19,428 22,481
[…]
M2 CPP 240 18
[…]
HE 4.5485 % 0.2954%
  • TRP– Total Received Packets; CPP– Correctly Predicted Packets; HIR— Human Intervention Required; HE— Human Efforts;
4.6 Comparison With Benchmark Studies
@@
Many studies validate their malicious traffic detection approaches in offline settings, often achieving significant results. However, these systems frequently experience performance degradation when deployed in real-time environments due to the dynamic nature of network traffic and the emergence of new attack patterns. The datasets used to train these models are typically static, which limits their effectiveness in real-world applications. There are no existing studies that evaluate their work on real-time online M-En traffic. Therefore, for a fair comparison, we implemented existing M-En studies on our newly collected real-time M-En dataset, deploying them according to the methodologies outlined in their respective published work. Since most existing studies conducted only offline testing, we replicated their models in the same offline setting to maintain consistency in evaluation. For instance, Rustam et al. [rustam2023securing] deployed their approach on an M-En dataset using the S-DATE. While the performance appeared strong after applying the S-DATE data balancing method, the results indicated potential data leakage issues, which could explain the observed high accuracy. Similarly, other studies [rustam2024malicious, rustam2024famtds] proposed optimization-based approaches to handle diverse traffic patterns in M-En networks, employing PSO and MFO, respectively. These studies introduced PSO-D-SEM and a Fully Automated Malicious Traffic Detection System (FAMTDS). Another study [rustam2024malicious1] utilized a dual-data trained LightGBM (DDT-LightGBM) model, achieving significant results in the M-En network within an offline testing framework. Zukaib et al. [zukaib2024meta] integrated federated learning and meta-learning to propose the meta-fed IDS, which was tested on the M-En dataset. However, their evaluation was limited to offline settings and did not incorporate real-time data collection or adaptation. Furthermore, we deployed modern GNN-based and transformer-based architectures to compare our approach with advanced approaches [10948513, 11199889]. In contrast, our study conducted both online and offline testing and was built on a real-time M-En dataset.
@@
In Table 19, we compare our approach, Multi-LF, to relevant M-En benchmark studies. Most existing approaches [rustam2023securing, rustam2024malicious, rustam2024famtds, rustam2024malicious1, zukaib2024meta] use static datasets and evaluate performance only in offline scenarios, reporting accuracy values between 0.94 and 0.991. While these systems can be effective in controlled settings, they often lack the mechanisms to continuously learn from new traffic patterns or adapt to real-time fluctuations in M-En networks. In contrast, Multi-LF is the only method evaluated in both offline and online modes, achieving near-optimal accuracy scores of 1.00 (offline) and 0.999 (online). These results highlight Multi-LF’s resilience to dynamic traffic behavior. By continuously retraining on fresh data and utilizing multi-level validation checks, Multi-LF maintains high accuracy even under unpredictable network conditions, offering a robust solution for real-time malicious traffic detection in M-En environments.
Table 19: Comparison With Existing Studies
@@
[…]
[rustam2024malicious] 2024 PSO-D-SEM $\checkmark$ $\times$ 0.978
[…]
[rustam2024malicious1] 2024 DDT-LightGBM $\checkmark$ $\times$ 0.980
[…]
[10948513] 2025 GTAE-IDS $\checkmark$ $\times$ 0.968
[…]
Our 2025 Multi-LF $\checkmark$ $\checkmark$ 1.00, 0.999
5 Discussion
In this study, we collect a dataset in the M-En and propose Multi-LF for malicious traffic detection, testing it in both offline and real-time scenarios. We deploy Multi-LF to reduce computational costs and enhance accuracy over time. We reduce the computational cost because the initial traffic is evaluated by the lightweight M1 model, which has a faster prediction time due to its lightweight architecture. Most of the traffic is filtered by M1, and only a small portion is forwarded to M2 and, if necessary, to a human expert. Additionally, we improve the approach’s performance by incorporating an extra layer of security using M2 and human involvement.
We achieve continuous improvement in model performance over time through continuous training with new data, enabling the model to efficiently adapt to evolving traffic patterns, as well as through our feature engineering strategy. We utilize general and statistical features, as they both significantly contributed to the model’s effectiveness, as illustrated in Figure 9. Statistical features, such as ConnectionErrors, DstPortEntropy, MostFreqPayloadSize, SourceEntropy, MostFreqPacketSizeFreq, FlowRate, PacketCount, PacketSizeVar, and AvgPayloadSize, played a critical role in detecting malicious traffic according to the importance score. For instance, a sudden spike in packet count from a specific IP within a short time window may signal an attack. Additionally, some general features, such as SYN and ACK flags, are also among the top contributors for malicious traffic detection. These statistical and general features together provided highly correlated input representations, as shown in Figure 9, thereby enhancing the overall performance of the framework.